OpenStudy (anonymous):

Find the particular solution for y''-4y'-5y= -6xe^-x

OpenStudy (anonymous):

r^2-4r-5 (r-5)(r+1) Ce^-t + C e^5t so no worry about linearity

OpenStudy (anonymous):

Can you show the process?

OpenStudy (anonymous):

sorry we do have to worry about linearity since your input function -6xe^-x has e^-x

OpenStudy (anonymous):

so we try (Ax^2+Bx+C)e^-x

OpenStudy (anonymous):

put it into your function

OpenStudy (anonymous):

I dont get it

OpenStudy (anonymous):

you don't get what?

OpenStudy (anonymous):

How to solve this

OpenStudy (anonymous):

we haven't solved it yet

OpenStudy (anonymous):

((Ax^2+Bx+C)e^-x)'' - 4((Ax^2+Bx+C)e^-x)' -5((Ax^2+Bx+C)e^-x)

OpenStudy (anonymous):

and you have to solve for A , B and C

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

when I simplify it , I get \[2 e^{-x} (A-3 B-6 A x)\]

OpenStudy (anonymous):

we set it equal to your input equation -6xe^-x

OpenStudy (anonymous):

-12 A=-6 A=1/2

OpenStudy (anonymous):

2A-6B=0 1-6B=0 B=1/6

OpenStudy (anonymous):

A x^2 + B x ) e^-x 1/2 x^2 +1/6 x) e^-x